OpenStudy (anonymous):

if secx+tanx/secx-tanx=5/3 then sinx is equal to.please have theta instead of x

OpenStudy (cp9454):

do the cross multiply

OpenStudy (cp9454):

\[3\sec \theta+3\tan \theta = 5\sec \theta -5\tan \theta\]

OpenStudy (cp9454):

\[8\tan \theta = 2 \sec \theta\]

OpenStudy (anonymous):

8cosx/sinx=2/cosx sinx=1/4

OpenStudy (cp9454):

\[\frac{ 4\sin \theta }{ \cos \theta } = \frac{ 1 }{ \cos \theta}\]

OpenStudy (cp9454):

u will get sin theta = 1/4

OpenStudy (anonymous):

please tell me how to solve prove like these identities.if i search in net i have only basics.i cant able to solve this kind of problems.if any shortcut is there means please help me